Following information were extracted from a Printing organization related to the sale of Articles in a particular year.

No of articles sold - 10,000

Selling price of an article - Rs. 30 per unit

Variable cost - Rs. 18 per unit

Fixed cost - Rs. 24,000 per year

Calculate the following.

(i) The profit of the business.

(ii) Break-even point in terms of money value and in units.

(iii) Margin of safety in terms of money value and in units.


1
Expert's answer
2021-07-29T18:18:02-0400

(i)                 

Profit = sales = fixed costs - variable costs

=300,000 - 180,000 - 24,000

=Rs 96,0000

 

(ii)

Contribution margin = selling price of the articles  – variable costs of the articles

= 30 – 18

= Rs 18

Break even in units = fixed costs of the articles /contribution margin

=24000/18

= 1,334 articles

Break even value = break even units x selling price of the articles

=1334 x 30

= Rs 40,020

 

 

 

(iii)

Margin of safety = (total sales – breakeven point sales)/ selling price of the articles

= (300,000 – 40,020)/30

= 8,666 articles

Margin of safety = 8,666 * 30

= Rs 259, 980

 

References

Sintha, L., 2020. Importance of Break-Even Analysis for the Micro, Small and Medium Enterprises. International Journal of Research-Granthaalayah, 8(6).